A 120.0 mL sample of 0.040 M Ca2+ is titrated with 0.040 M EDTA at pH 9.00. The value of log⁡Kf for the Ca2+−EDTA complex is 10.65 and the fraction of free EDTA in the Y4− form, αY4−, is 0.041 at pH 9.00.

What is Kf′, the conditional formation constant, for Ca2+ at pH 9.00?

Kf′=

What is the equivalence point volume, Ve, in milliliters?

Ve=    mL

Calculate the concentration of Ca2+ at V=12Ve.

[Ca2+]= M

Calculate the concentration of Ca2+ at V=Ve.

[Ca2+]= M

Calculate the concentration of Ca2+ at V=1.1Ve.

[Ca2+]= M
Answer #1

Consider a complex formation reaction: Ca 2+ + EDTA phpHNPpXX.png [CaY] 2- ------------------->(1)

For above reaction, conditional formation constant is K' f = (phpOFEv2K.pngY 4-) K f

Therefore, K' f = 0.041 x 10 10.65 = 1.83 x 10 9

Calculation of volume of EDTA required to reach equivalence point

From above reaction (1), we have 1 mol Ca 2+\equiv 1 mol  EDTA

Therefore, M ca 2+ \times V Ca2+ = M EDTA\times V EDTA

phplu1Wsc.png V EDTA = M ca 2+ \times V Ca2+ / M EDTA =

V EDTA = 0.040 M \times 120.0 ml / 0.040 M = 120.0 ml

phpgmYUxF.png Equivalence point volume = 120.0 ml

[Ca 2+] at V = 1/2 Ve

We know that, Molarity = No .of moles of solute / volume of solution in L

Therefore, No .of moles of solute = Molarity \times volume of solution in L

No. of moles of Ca 2+ present initially in the solution = 0.040 mol / L \times 0.1200 L = 0.0048 mol

No. of moles of EDTA added in the solution = 0.040 mol / L \times 0.0600 L = 0.0024 mol

Excess moles of Ca 2+ left in the solution after reaction with EDTA = 0.0048 - 0.0024 = 0.0024 mol

Volume of solution at this stage = Volume of Ca 2+ solution + volume of EDTA solution = 120.0 + 60.00 ml = 180.0 ml

[Ca 2+ ] = 0.0024 mol / 0.1800 L = 0.0133 M

[Ca 2+] at V = Ve

At the equivalence point, all the calcium ions are consumed by added EDTA. Hence, at this stage concentration of Ca 2+ will be due to dissociation of [CaY] 2-.

Moles of [CaY] 2- = moles of EDTA added to the solution

No. of moles of EDTA added in the solution = 0.040 mol / L \times 0.1200 L = 0.0048 mol

Volume of solution at this stage = Volume of Ca 2+ solution + volume of EDTA solution = 120.0 + 120.0 ml = 240.0 ml

[CaY] 2-. = 0.0048 mol / 0.2400 L = 0.0200 M

Let’s use ICE table.

Concentration Ca 2+ EDTA [CaY] 2-
Initial 0.0200
Change +X +X -X
Equilibrium X X 0.0200 - X

Therefore, K' f = [CaY2-] / [Ca 2+] [EDTA] = 0.0200 -X / (X)(X) = 1.83\times10 9

0.0200 -X / (X)(X) = 1.83 \times 10 9

phpS0fX8v.png 0.0200  -X / X 2 = 1.83 \times 10 9

At equilibrium X is negligible as compared 0.0200. Then we can write 0.0200 / x 2 = 1.83 x 10 9

phpmJXRa5.png X 2 = 0.0200 / 1.83 \times 10 9 =1.093 \times 10 -11

php2kK1gu.png X = 3.306 x 10 - 6 M

phpA7fuBx.png [Ca 2+] = 3.306 x 10 - 6 M

[Ca 2+] at 1.1 Ve (1.1 x 120 = 132 ml)

After equivalence point, there is excess EDTA its concentration is calculated as shown below

No. of moles of Ca 2+ present initially in the solution = 0.040 mol / L \times 0.1200 L = 0.0048 mol

No. of moles of EDTA added in the solution = 0.040 mol / L \times 0.1320 L = 0.00528 mol

Excess moles of EDTA = 0.00528 - 0.0048 = 0.00048 mol

No. of moles [CaY 2-] produced = No. of moles of Ca 2+ present initially in the solution = 0.0048 mol

Volume of solution at this stage = Volume of Ca 2+ solution + volume of EDTA solution = 120.0 + 132.0 ml = 252.0 ml

[EDTA] = 0.00048 mol / 0.2520 L = 1.905 \times 10 -03 M

[CaY 2-] = 0.0048 mol / 0.2520 L = 0.01905 M

We have, K' f = [CaY2-] / [Ca 2+] [EDTA] = 1.83 \times 10 9

php6ZGC7N.png (0.01905 ) / [Ca 2+] (1.905 \times 10 -03) = 1.83 \times 10 9

phpDlxZsV.png [Ca 2+] (1.905 \times 10 -03) = (0.0195) / 1.83\times 10 9 = 1.06 \times 10 -11

phplndnXz.png  [Ca 2+] = 1.06 x 10 -11/ (1.905 \times 10 -03) = 5.6\times 10 -09 M

[Ca 2+] = 5.6 x 10 -09 M